### v2

# Particle migration in areas of constricted flow

Abstract
Cardiovascular diseases are a leading cause of death globally. Among them, some are linked to stenosis, which is an abnormal narrowing of blood vessels, as well as other factors. Smart drug delivery systems based on micro- and nanoparticles are a promising method to offer non/minimal-invasive therapeutic mechanisms. Here we investigate the propensity of particles with different shapes and sizes to drift laterally (marginate) towards an occlusion area in a two-dimensional (2D) parallel plate laminar flow using the Lattice-Boltzmann method (LBM). To verify the outcomes on both sides of the stenosis, a probability of adhesion to the borders was calculated. Analysis was done on the impact of wall-shear stress on both sides of the stenosis. Our results show that rectangular particles migrate in larger amounts and earlier than circular ones.
keywords:
Lattice-Boltzmann method , Particle dynamics , Blood flow , Constricted flow
journal: Communications in Nonlinear Science and Numerical Simulation
[…]
1 Introduction
Particle migration in flows refers to the transport of small particles or droplets within a fluid as it flows through a system. This can occur in a variety of contexts, such as in rivers, pipelines, and even in the human body. Particle migration can be affected by several factors, including the size and shape of the particles [1], the properties of the fluid, and the presence of external forces such as gravity or magnetism. Understanding particle migration is important in a wide range of fields, including chemical engineering, environmental science, and medicine [2].
There has been a great deal of research on particle migration in flows in recent years, and significant progress has been made in understanding the underlying mechanisms and developing predictive models [3]. One area of active research is in developing methods to control or manipulate particle migration in flows, such as using external fields or by altering the properties of the fluid or particles. Another focus is on understanding the impact of particle migration on the transport of heat, mass, and momentum in a system, as well as on the overall stability and performance of the flow [4, 5]. Additionally, much work has been carried out on developing experimental techniques to measure and characterize particle migration in flows. Notable methods from previous decades include the use of optical scanning devices coupled with electronic counting circuits [6], and setting large reservoirs to pump red cell-plasma suspensions through capillaries of different sizes [7]. More recently, advanced techniques such as laser-based imaging or tracking systems [8] have emerged as powerful tools to study this phenomenon. Interaction of particles with solid and elastic obstacles in the flow may lead to changes in lateral migration.
Nanoparticles, particularly spheres, are typically proposed as carriers for vascular-targeted drug delivery due to their relative ease of fabrication and their capacity to navigate the circulatory system with minimal risk of vessel occlusion [9]. However, injected targeted therapeutics have often led to suboptimal and adverse systemic effects due to several factors, including early lateral migration and an overall inability to reach the designated target [10]. Some investigations have shown that efficacy of carriers for drug delivery depends on their ability to marginate (localize and adhere) to the vessel wall [11]. In this regard, several studies suggest that, for particles circulating in a flow, the strategy of using larger particles to marginate could be more effective than using small ones [12]. In fact, particles with submicrometer diameters have a lower capacity to bind the endothelium possibly because they are entrapped within the main flow [13]. Furthermore, recent research shows the benefits of using non-spherical particles that can favor margination and interaction with the endothelium [14]. Particle size, shape, core-shell structure and stiffness play important roles in influencing particle behavior in blood flow in terms of cell-particle and vessel-particle interactions, transport and lateral margination. Computational models [15] have suggested that spherical particles tend to follow streamlines, while non-spherical particles (discs, rods…) can suffer significant lateral drift, even in the absence of external forces, which favors migration [16, 17, 18, 19].
Several types of shear stress-sensitive drug vehicles are under research, in which drug delivery would be triggered by shear stress changes [20, 21, 22]. Recently, there has been an increased interest in shear-activated particle clusters for the treatment of stenosis [23]. However, to the best of our knowledge, there are no drug-carriers guided to the vascular lesion (by the flow characteristics and the change in wall shear stress) which enable a localized internalization in the diseased cells/tissues while providing a sustained release of the therapeutic cargo.
In recent years, Lattice-Boltzmann methods (LBMs) [24] have evolved enough to compete with classical tools in solving complex fluid flows problems, and be able to simulate particulate flows. Rooted in the mesoscale, the LBM models the fluid as a collection of particle distribution functions, and simulates their evolution over time. In this work, LBM were used to study the lateral migration of particles of different shapes and sizes in a 2D stenotic flow. In the current study, a Lagrangian particle tracking scheme was combined with the LBM scheme.
Regarding the characterization of the blood, whole blood is a complex suspension made of a liquid phase, the plasma, and the cellular elements suspended within it, including red blood cells (RBCs), white blood cells, and platelets. Plasma itself behaves as a Newtonian fluid [25]; in contrast, whole blood exhibits non-Newtonian behaviour primarily due to the presence of RBCs. Nevertheless, for many applications blood can be considered a Newtonian fluid, provided that the wall shear stress observed in the regions of study are not too low [26, 27]. In such cases, the rheological differences in blood behavior are negligible and the Newtonian assumption holds true. Since this study focuses exclusively on the plasma, the use of a Newtonian model is therefore justified.
2 Numerical method
To study the motion of a particle cluster in a 2D constricted channel, the Lattice-Boltzmann method (LBM)[28, 24, 29] is adopted. The dynamics of the Newtonian fluid flow with an external force can be described by a single relaxation time via the Bhatnagar-Gross-Krook (BGK) equation [30],
$f_{i}({\bf{r}}+{\bf{e}}_{i}\Delta t,t+\Delta t)-f_{i}({\bf{r}},t)$ $=$ $\frac{\Delta t}{\tau}\left[f_{i}^{eq}({\bf{r}},t)-f_{i}({\bf{r}},t)\right]$ (1)
$+\Delta tS_{i}$
where $f_{i}(\bf{r},t)$ is the distribution function at position $\bf{r}$ and time $t$, $f_{i}^{eq}(\bf{r},t)$ is the equilibrium distribution function [31], $\tau$ is the dimensionless relaxation time, and $\Delta t$ is the time step. In the present study, the D2Q9 (2-dimensional and 9-velocity) model [24] is adopted, where the equilibrium distribution function is defined as,
$f_{i}^{eq}=\omega_{i}\rho\left[1+\frac{{\bf{c}}_{i}\cdot\bf{u}}{c_{s}^{2}}+\frac{\left({\bf{c}}_{i}\cdot\bf{u}\right)^{2}}{2c_{s}^{4}}-\frac{u^{2}}{2c_{s}^{2}}\right]$ (2)
Here $\rho$ and $\bf{u}$ are the fluid density and velocity, respectively; $\omega_{i}$ represents the weighting parameter of the model, where $\omega_{0}=4/9$, $\omega_{1-4}=1/9$ and $\omega_{5-8}=1/36$; ${\bf{c}}_{i}$ is the discrete velocity of the model, with $c_{0}=(0,0)$, $c_{1-4}=c(\cos[(i-1)\pi/2],\sin[(i-1)\pi/2])$, $c_{5-8}=2c(\cos[(2i-1)\pi/4],\sin[(2i-1)\pi/4])$; $c=\Delta x/\Delta t$ is the lattice speed, which was assumed to be one since $\Delta x=\Delta t=1$; $\Delta x$ is the lattice spacing; and $c_{s}=c/\sqrt{3}$ is the sound speed of the model. The last term in Eq. (1) represents the source term, which takes into account the force density driving the fluid as well as the forces exerted by particles, for which we use the Guo forcing scheme [32]. Based on this scheme, a body force density $\bf{F}$ is related to $F_{i}$ as,
$S_{i}$ $=$ $\left(1-\frac{\Delta t}{2\tau}\right)F_{i}$ (3)
$F_{i}$ $=$ $\omega_{i}\left(\frac{\bf{c_{i}}-\bf{u}}{c_{s}^{2}}+\frac{\bf{c_{i}}\cdot\bf{u}}{c_{s}^{4}}\bf{c_{i}}\right)\bf{F}$ (4)
The local fluid density, $\rho$, and velocity, $\bf{u}$, at the lattice node are given by $\rho=\sum_{i}f_{i}$ and $\rho{\bf{u}}=\sum_{i}f_{i}{\bf{c_{i}}}+\Delta t{\bf{F}}/2$. $\bf{F}$ is the sum of the forces acting on the fluid. The LBM for a single-phase flow recovers the Navier-Stokes equation in the limit of small Knudsen number for incompressible fluids, with the fluid kinematic viscosity being $\nu=c_{s}^{2}(\tau-0.5)$ (in lattice units). The kinematic viscosity cannot be less than zero; therefore, the limitation $\tau>0.5$ must be satisfied.
The particle–fluid interaction was implemented using a force-coupling immersed boundary lattice Boltzmann method (IB-LBM), conceptually following the formulation of Safdari and Kim [33]. In this approach, the particle surface is discretized into uniformly distributed boundary nodes, whose velocities are prescribed by the rigid-body motion of the particle. At each boundary node, the surrounding fluid velocity is interpolated and compared to the particle’s velocity to evaluate the hydrodynamic forces and torque contributions. This approach differs from the dissipative-coupling method proposed by Ahlrichs and Dünweg [34], as no friction parameter is required, and from the momentum exchange method of Niu et al. [35] or the classic IBM [36], since our implementation explicitly evaluates drag and lift contributions, which are subsequently spread back to the fluid through Eq. (4), rather than via a discrete delta function.
The particles trajectories and velocities can be calculated from a Lagrangian point of view by integrating,
$m_{p}\frac{d{\bf{v}}_{p}}{dt}$ $=$ $\bf{F_{D}}+\bf{F_{L}}$ (5)
$I_{p}\frac{d\Omega_{p}}{dt}$ $=$ $\bf{T}$ (6)
where $m_{p}$ and $I_{p}$ denote the mass and moment of inertia of the $p$ particle, ${\bf{v}}_{p}$ and $\Omega_{p}$ are the translational and rotational velocities of each particle, $\bf{F_{D}}$ represents the drag force, $\bf{F_{L}}$ is the lift force, and $\bf{T}$ the particle torque. Note that the effect of forces like virtual mass, Basset history and Magnus effect can be considered negligible. Other forces (Brownian, gravity, etc) acting on the particles are not considered in this paper. Particles’ boundary are discretized in $M$ nodes as seen in Fig. 1, whose velocities ${\bf{v}}_{b}$ are given by,
${\bf{v}}_{b}={\bf{v}}_{p}+\Omega_{p}\times({\bf{r}}_{b}-{\bf{r}}_{p})$ (7)
with ${\bf{r}}_{b}$ and ${\bf{r}}_{p}$ being the boundary node and particle positions, respectively.
Figure 1: Representation of a single discretized circular particle (red dots) moving through the fluid lattice (white dots). Each particle boundary node is assigned a velocity $\bm{v_{b}}$. The corresponding fluid velocity at the same location, $\bm{u_{b}}$, is obtained via a four–neighbor interpolation, illustrated by the gray rectangle.
Finally, the drag and lift forces and the torque are given as,
$\bf{F_{D}}$ $=$ $C_{D}\sum_{b}({\bf{u}}_{b}-{\bf{v}}_{p})$ (8)
$\bf{F_{L}}$ $=$ $C_{L}\sum_{b}\left(({\bf{u}}_{b}-{\bf{v}}_{p})\times\left[{\bf{\nabla}}\times({\bf{u}}_{b}-{\bf{v}}_{p})\right]\right)$ (9)
$\bf{T}$ $=$ $C_{T}\sum_{b}\left(({\bf{r}}_{b}-{\bf{r}}_{p})\times({\bf{u}}_{b}-{\bf{v}}_{p})\right)$ (10)
where $C_{D}$, $C_{L}$ and $C_{T}$ denote the drag, lift, and torque coefficients, respectively, and ${\bf{u}}_{b}={\bf{u}}({\bf{r}}_{b})$ is the fluid velocity at the boundary nodes. The summation runs over all $M$ boundary nodes. For the lift equation (9) we considered the Saffman force [37] with $C_{L}=0.5$ for the sake of simplicity, although more complex models could be used [38]. The drag coefficient is taken as a function of the particle Reynolds number, following the Schiller–Naumann correlation [39] (see also [40]). We further assume that the torque coefficient equals the drag coefficient, i.e $C_{T}=C_{D}$.
Then, the position and velocity of each particle is updated by a forward explicit Euler scheme:
${\bf{r}}_{p}(t+\Delta t)={\bf{r}}_{p}(t)+\Delta t{\bf{v}}_{p}$ (11)
${\bf{v}}_{p}(t+\Delta t)={\bf{v}}_{p}(t)+\Delta t{\bf{F}}$ (12)
with $\bf{F}=\bf{F_{D}}+\bf{F_{L}}$. The accuracy of this coupling procedure, as well as the overall implementation, has been validated against standard benchmarks, which are reported in the Supplementary Material.
Figure 2: Schematic of the 2D setup used in the simulations.
Fig. 2 shows the 2D rectangular domain ($H\times L=400\times 4000$ cells) considered (see A for a grid convergence study). The constriction zone or stenosis is simulated as a Gaussian function of height $h_{0}$ and variance $h_{0}^{2}$ centered at $L/2$. A Poiseuille flow with a maximum velocity $u_{max}$ at $H/2$ was considered at the inlet and an outflow boundary condition was imposed for the outlet, both achieved through the use of Zou/He type boundary conditions on the mesoscopic level [41]. The bounce-back scheme was implemented on channel top and bottom walls—including the curved Gaussian regions—to simulate the no-slip condition on the fluid, capturing the wall and inertial effects on particle-fluid motion, as discussed by Basagaoglu et al. (2008)[42].
Particles had initially the same velocity as the flow at the release location. In this paper, we focus specifically on circular and rectangular rigid particles. $N$ particles were released randomly far enough from the constriction zone and at $H/4$ cells far from the lateral boundaries (gray area in Fig. 2). Care was put in making sure there is no overlapping between them. Rectangles were set initially at random orientations. For a circular particle with radius $R$ rotating around its center of mass, the angular moment of inertia is $I_{p}=m_{p}R^{2}/2$, while for a rectangular particle with sides $l_{1}$ and $l_{2}$, $l_{2}\geq l_{1}$, $I_{p}=m_{p}(l_{1}^{2}+l_{2}^{2})/12$. We consider elastic collisions between two particles, or a particle and a wall. For the collision between circular particles we suppose a hard sphere collision, whereas for the collision between two rectangles we follow Qin et al. (2020) [43]. For the last case, the collision model takes into account whether the collision involves two vertices, a vertex and an edge or two edges. In this paper, we do not consider collisions between circular and rectangular particles. To compare results between both types of particles, we define the equivalent radius for the rectangle particles as $R_{equiv}=l_{1}(1+\gamma^{2})^{1/2}/2$ where the aspect ratio is defined as $\gamma=l_{2}/l_{1}$. The fluid is initialized with a density $\rho=1$, velocity $u_{max}=0.04$, Reynolds number $Re=Hu_{max}/\nu$ and a relaxation time $\tau=0.62$, all given in lattice units. Henceforth, all results presented in this work are expressed in lattice units.
3 Results
3.1 Flow trajectories of particles
The LBM was used to simulate the migration trajectories of particles as a function of their size and shape. Simulation results in Fig. 3 unveiled two distinct shape dependent particle behaviors in a constricted channel: (i) both circular and rectangular particles migrate towards the lateral boundaries; however, the rectangles do so before the circles. Additionally, (ii) rectangular particles migrate to both sides of the stenosis and in greater quantity than circles. This finding aligns with previous studies focusing on the behaviour of particles with different aspect ratios ranging from spherical to discoidal [44]. Note that both types of particles create a wake, in which others are caught and therefore experience less drag, resulting in slower motion and eventual clustering. This is less pronounced for rectangles as they displayed non-zero cumulative angular rotations favoring their dispersion, unlike the circular particles.
Particles roll along the walls of the geometry. Their probability of adhering to the boundaries will be considered later. One may find a small overlapping area of touching particles, especially for rectangles. This is expected from the numerical method because of the smooth transition of the object boundaries. The size of the overlapping area was reduced to its minimum size adjusting the relaxation time $\tau$.
Figure 3: Flow velocities and particles positions for different shapes; (a-b) circular, $R=10$ and (c-d) rectangular, $R_{equiv}=15.8$ ($l_{1}=10,\gamma=3$) at different times crossing the constriction zone. In every case, the initial positions of the particles were the same. The color bar shows the values of flow velocity. Set of parameters: $Re=400$ and $N=40$.
To investigate the dispersion of particles with different aspect ratios we calculate the mean square displacement (MSD) as a function of time,
$MSD(t)=\langle(\bm{r}_{k}(t)-\bm{r}_{k}(0))^{2}\rangle$ (13)
where $\bm{r}_{k}$ is the position of the k-th particle.
Fig. 4 shows the increase rate of the mean square displacement for different particle diameters. The rate of increase is higher for a smaller $R_{equiv}$, which is caused by the smaller particles having a faster momentum response time with respect to the fluid flow. Increasing the particles aspect ratio increases interparticle interactions and particle dispersion with time.
Figure 4: Increasing rate of the MSD as a function of the equivalent radius for circles and rectangles. The rate was calculated as the slope of a log-log fitting of the MSD(t). Circle and square dots represent circular and rectangular particles, respectively. Percentage occlusion: 25%, and rest of parameters as in Fig. 3.
Particle migration to the edges depends on their size and shape as their response to drag and lift forces is different. Particles are considered to have migrated to the edges if the distance from any particle node to the boundaries is less than 5 cells and they are in the intervals [L/2L/8,L/2][L/2-L/8,L/2] before and $[L/2,L/2+L/8]$ after the stenosis. Fig. 5 shows the ratio of migrated particles $N_{migr}/N$ for circles and rectangles as a function of $R_{equiv}$ before (a) and after (b) the stenosis. Note that for both occlusion percentages, $N_{migr}$ is larger for the rectangular particles than for the circular ones. Small circular particles hardly see their trajectory modified and are instead confined in the center-region of the channel. Before the stenosis, $N_{migr}$ increases with $R_{equiv}$ in all cases, while the number of circles reaching the boundaries after the stenosis decreases as their size increases. Large particles have larger residence times, Fig. 5(c), as they were less frequently trapped by the recirculating vortices. Residence time increases with the size of the occlusion zone and the recirculating vortices.
Figure 5: Ratio of particles migrated to the lateral boundaries and mean residence time at the boundaries before (a,c) and after (b,d) the stenosis as a function of $R_{equiv}$. Blue-solid lines correspond to an occlusion of 25% and red-dashed lines to 50%. Circles and squares dots are for circles and rectangles particles, respectively. Set of parameters as in Fig. 3.
Once particles cross the stenosis, Fig. 5(b), circles are dragged by the central flow, where the velocity is at a maximum ($N_{migr}$ decreases with increasing $R_{equiv}$). For the rectangles, lift forces increase and they have a higher probability of migrating to the boundaries. Residence times do not change appreciably with $R_{equiv}$, but, in contrast to the previous case, decrease as the percentage of occlusion increases, Fig. 5(d). Very large particles can collapse/stop the flow at the stenosis by stacking [45]. It should be noted that in three-dimensional (3D) simulations, although particle accumulation may occur, the fluid velocity would not drop to zero, as there is space for the fluid to flow around the particles even if they partially obstruct the stenosis throat[46].
The role of the initial positioning of the particles was also investigated, although the results are not shown here. Particles injected far from the stenosis are more likely to be dragged by the central flow, with this phenomenon being more probable for circular particles than for rectangular ones.
The influence of the occlusion percentage on particle migration is shown in Fig. 6. For both types of particles with similar values of $R_{equiv}$, the number of particles that migrate to the boundaries increases with an increasing occlusion; in fact, this trend is not exclusive to artificial particles, as it was also observed by Yazdani and Karniadakis in a similar study on platelet margination in the presence of RBCs [47]. Wall-shear stress upstream of the stenosis increases with $h_{0}$, favoring migration of particles to the boundaries; conversely, post-stenosis wall-shear stress diminishes due to the presence of a recirculation zone, thus reducing the rate of particle migration, especially for circular particles, which are more prone to be dragged by the flow. The driving forces for margination are significantly altered by the constriction.
Figure 6: Ratio of particles migrated to the lateral boundaries before (a) and after (b) the stenosis as a function of the occlusion percentage. Blue line corresponds to circles with $R=10$ and red line to rectangles with $R_{equiv}=9.5$ ($l_{1}=6,\gamma=3$). Set of parameters as in Fig. 3.
Fig. 7 illustrates the influence of the number of released circular particles $N$ and the Reynolds number on the ratio of particles that migrated to the boundaries. All Reynolds numbers considered lie within the physiologically relevant range for standard arteries. Only results containing circular particles are displayed; earlier results demonstrating rectangular particles indicate that $N_{migr}$ is greater than that of circular particles. As expected, $N_{migr}$ increases with $N$ before and after the stenosis, but increasing $N$ leads to larger slow flow at the particles wake and clogging at the constriction, thus reducing the ratio of particles that migrate to the lateral boundaries with respect to the number of emitted particles. This is more evident after the stenosis. On the other hand, increasing the Reynolds number by increasing $u_{max}$ leads to a linear increase of $N_{migr}$, as the drag and lift forces are proportional to $Re$.
Figure 7: Ratio of circular particles migrated to the lateral boundaries before (blue-solid line) and after (red-dashed line) the stenosis as a function of the number of released particles (a) and the Reynolds number (b). Percentage of occlusion 25%, $R=10$, $Re=400$ (a) and $N=40$ (b).
3.2 Probability of adhesion
Particle migration occurs naturally due to flow conditions, being faster and in greater quantity for non circular shapes. However, particles were not intended to adhere to the boundaries so they may roll down the walls and thus modify the number of particles migrating to either side of the stenosis.
There are numerous models for calculating the general adhesion strength [48, 49]. In this paper we used a simplified version of the bioadhesion model by Decuzzi and Ferrari [50], in which the adhesion strength directly correlates with the higher chance of successful bonds of particles to the boundaries, $P$. In the original model, the adhesive strength of the particle arises from the selective binding between receptors on the target surface and complementary ligands displayed on the particle surface. and the strength of adhesion is determined by geometric factors (particle volume and aspect ratio), biophysical parameters (equilibrium separation distance, maximum bonding distance, wall shear stress relative to receptor density, and the densities of ligands and receptors), and biochemical parameters (bond characteristic length and affinity constant of the ligand–receptor pair).
To use this model we assume the rectangles to be oval-like particles with aspect ratio $\gamma$, and we adopt the same threshold of 5 cells as in the margination calculation (Section 3.1), beyond which adhesion is not considered possible. Thus, for particles of circular shape,
$P$ $=$ $\pi r_{0}^{2}\exp\left[-2.4\cdot 10^{-4}fR_{equiv}\mu_{s}/r_{0}^{2}\right]$ (14)
$r_{0}^{2}$ $=$ $R_{equiv}^{2}\left[1-\left(1-0.05\gamma/R_{equiv}\right)^{2}\right]$ (15)
$f$ $=$ $6\left(R_{equiv}\gamma^{-1}+1\right)F^{s}+8R_{equiv}^{2}T^{s}/r_{0}$ (16)
where $F^{s}$ and $T^{s}$ are coefficients that hinge upon the aspect ratio of particles $\gamma$ [50], and $\mu_{s}$ is the wall-shear stress. Eq. 14 shows that, in this formulation, adhesion strength is primarily governed by the geometric properties of the particle, namely its radius $R_{equiv}$ and the aspect ratio $\gamma$.
The variation of the adhesion probability $P$ with the radius $R_{equiv}$ for three different wall shear stress regions is shown in Fig. 8(a). These regions consist of three sectors: a shear acceleration zone at the proximal end of stenosis (circles and blue solid line), a peak shear zone just before the apex of stenosis (squares and red dash-dotted line), and a shear deceleration zone at the distal end of stenosis within the recirculation area (crosses and dashed line). Zones with smaller wall shear stress show a larger adhesive probability [50]. For any of the sectors, $P$ reaches a maximum value for some intermediate optimal value of the radius and then decreases. Results not shown for rectangular particles lead to an overall increase of the adhesive probability $P$ as the aspect ratio increases. All these observations are consistent with [50].
Figure 8: Adhesive probability for three different wall shear stress regions (a) and ratio of migrated (red upper line) and adhered (blue lower line) circular particles before the stenosis (b) as a function of $R_{equiv}$. Set of parameters: $Re=400$, $N=100$ and percentage of occlusion 50%.
As mentioned in the previous section, an increasing number of particles migrate to the boundaries upstream of the stenosis as their radii increases, as shown in Fig. 8(b). However, the number of particles adhered to the walls decreases with increasing radius while $P$ decreases (Fig. 8(a)). The optimal radius for which the number of adhered particles is maximum was not observed. That is most likely due to the minuscule probability differences that are seen for small values of $R_{equiv}$. Post-stenosis, fewer circular particles migrate and adhere to the boundaries, although $P$ increases due to a shear stress reduction. The flow acceleration at the apex of stenosis drags the particles and prevents their approach to the edges.
Adhesion of particles to the lateral boundaries does not happen continuously in time. This behavior can be described statistically by the waiting time distribution $q(\tau_{w})\propto\tau_{w}^{-\beta}$. The time intervals between two particles adhesion are known as waiting times $\tau_{w}$. The exact values of the exponent can be influenced by the finite length of the integration time or the number of particles, which underestimates the occurrence of long waiting times. Results show that the exponent $\beta$ is negligibly affected by $R_{equiv}$[50]. Fig. 9 shows the initial period of time $\tau_{w_{0}}$ for the first particle to adhere to the boundaries. Larger circles adhere faster than shorter ones because they migrate earlier and in larger quantities, but as previously said, they adhere less.
Figure 9: Initial period of time for circular particles adhesion $\tau_{w_{0}}$ to the lateral boundaries as a function of $R_{equiv}$. Dashed line is shown as a visual guide. Set of parameters: $Re=400$, $N=100$ and percentage of occlusion 50%.
4 Conclusions
The particle-particle and particle-walls interactions in a constricted channel have been studied using the Lattice-Boltzmann method, where the effects of particles size and shape, as well as their probability of adhesion to the boundaries, were investigated.
Rectangular-shaped particles migrate more than circular ones to the edges on both sides of the constricted zone. They also exhibit greater sensitivity to changes in fluid flow, having a faster momentum response than circular particles. However, in both cases, particle size remains the dominant factor, with smaller particles reacting more rapidly. Strikingly, the use of medium-sized rectangular particles, rather than large circular particles, resulted in enhanced entrapment on both sides of the stenosis. This could be used to increase the overall efficacy of carriers for drug delivery without the need for very big and/or a large number of particles that could occlude the vessel. Particle size also influences the mean residence time at the boundaries, increasing with the radius and the occlusion percentage before the stenosis, but remaining mostly constant with respect to the radius and inversely proportional to the occlusion percentage after the stenosis. Due to the presence of recirculating flow at the stenosis wake, particle migration was found to increase with increasing occlusion, especially for rectangles following the stenosis.
The adhesion strength favors early migration of particles to the boundaries on both sides of the constricted zone. Regions upstream of the stenosis with large wall-shear stress values have a higher probability of particle adhesion. Future research should consider the potential for particle aggregation leading to the formation of piles at the boundaries, which could result in particle clusters that significantly alter the flow velocity at the constriction. Results presented in the present work may aid for optimal design of intravascular injectable carriers near a stenosis.
It should be noted that extending the present analysis to three dimensions could alter the quantitative outcomes. In such geometries, secondary flows, radial shear gradients, and out-of-plane particle motions can modify migration dynamics compared to the 2D case presented here. Blockage of blood flow in the constricted region due to particle clogging is unlikely in 3D vessels, reducing the likelihood of complete obstruction. In addition, particle orientation and contact area with the vessel wall are more complex in 3D, potentially influencing adhesion probabilities. While the qualitative trends reported here—such as the dependence of margination and adhesion on particle size and aspect ratio—are expected to remain valid, the precise migration rates and adhesion efficiencies may differ. A full three-dimensional study would therefore be an important direction for future work.
CRediT authorship contribution statement
Raquel Dapena-García: Methodology, Conceptualization, Validation, Formal analysis, Writing - Original Draft. Vicente Pérez-Muñuzuri: Conceptualization, Writing - Review &\& Editing, Supervision, Funding acquisition, Project administration.
Declaration of competing interest
The authors have no conflicts to disclose.
Data availability
Source code for the LBM simulations is available from the authors upon reasonable request.
Acknowledgements
This research has been supported by the Xunta de Galicia (Grant No. 2025-PG025) and Ministerio de Ciencia e Innovación (Grant No. PID2022-141626NB-I00).
Appendix A Grid convergence details
The numerical method was evaluated for four distinct grids ($H=200$, $400$, $600$, $800$ and $L=10\times H$) for circular particles moving through a stenotic two-dimensional channel with two percentages of occlusion values at $Re=400$ and $N=100$. To that end, the time-averaged particle Reynolds number was calculated:
$Re_{p}>=\frac{1}{NT}\sum_{t,k}\frac{2R{\bf{v}}_{k}}{\nu}$ (17)
with ${\bf{v}}_{k}$ being the k-th particle velocity at time $t$. Particle radius ($R=5$, $10$, $15$, 2020), stenosis height $h_{0}$, and initial fluid velocity $u_{max}$ were renormalized appropriately as $H\times L$ was varied, and to keep the Reynolds number $Re=400$ constant with $\nu=0.04$, which was widely used in the results shown above. The number of time steps $T$ in each simulation was increased by a factor of 4 to simulate the same physical time in each case. $Re_{p}>$ calculated for the three larger grids are within $\approx 4\%$ error as shown in Table 1, with the discrepancies being due to the random initial particles positions, and low number of particles. As mean particle velocities in the central flow increase with occlusion, $Re_{p}>$ also does. Ultimately, in our simulations, to be computationally efficient, we have chosen the medium grid ($H\times L=400\times 4000$ cells).
$H\times L$ $Re_{p}>(25\%) <$$Re_{p}>(50\%)$
$200\times 2000$ 0.21 0.28
$400\times 4000$ 0.12 0.14
$600\times 6000$ 0.11 0.13
$800\times 8000$ 0.12 0.13
Table 1: Mean particle Reynolds number for two occlusion values (25% and 50%) and four grid sizes $H\times L$.